= Solve the following equations 2 1 a) 11_1 + 3x 4 6x 3 b) V nr3, solve for r. c) 25cos20 – 2 = 0 d) V7x + 2 + x = 6 3
Answer #1

르 + 31 ㅗ 4. ㅛㅗ 아. 3 - + L 3 - 4 ㅛ 2 6 3가 1-4 E + (2) 구 6L 크 62 2 VE 93 = -v V2 47 T- 이25cose -2 =0 ascost = 2 COSO 25 Coso 2 3- to ( ) 「+2+=6 5 6X both sides squaring j(+232 - (E-92 고 36+기를 120 은 190 434=O 기르 1